OpenStudy (anonymous):

if y=u+2e^u and u=1+lnx, find dy/dx when x=(1/e), please help i got as far as plugging everything in and getting y=1+ln(1/e)+2e^(1+ln(1/e))

OpenStudy (baru):

use chain rule \[\frac{ dy }{dx }=\frac{dy}{du}\times \frac{du}{dx}\]

OpenStudy (anonymous):

im not sure how to do that in this case

OpenStudy (anonymous):

oh wait, please hold on one moment

OpenStudy (baru):

ok

OpenStudy (baru):

like I showed in the formula differentiate y with respect to u differentiate u with respect to x multiply the two, you get dy/dx

OpenStudy (anonymous):

would you recommend differenciating every value first?

OpenStudy (anonymous):

for example if we know x=1/e then we use quotient rule

OpenStudy (baru):

noo, x is a variable, you can substitute x=1/e only after you have found dy/dx

OpenStudy (baru):

remember 'e' is a constant

OpenStudy (baru):

not another variable

OpenStudy (anonymous):

so we begin by finding the derivative of y=u+2e^u?

OpenStudy (baru):

yes, with respect to 'u'

OpenStudy (anonymous):

that means we plug in u already

OpenStudy (baru):

no, consider 'u' as a variable

OpenStudy (baru):

we are not plugging in anything

OpenStudy (anonymous):

ok, so when i differentiated y=u+=2e^u i got u+2(lne)(e^u)(u')

OpenStudy (anonymous):

and (lne)=1

OpenStudy (baru):

y=u + \(2e^u\) \[\frac{dy}{du}=1 + 2e^u\]

OpenStudy (anonymous):

so we don't have to find the derivative of 2e^u?

OpenStudy (baru):

according to the rules of differentiation\[\frac{ d(e^x) }{ dx }=e^x\]

OpenStudy (anonymous):

ok so what would the next step be?

OpenStudy (anonymous):

do we plug in u now?

OpenStudy (baru):

no, now u= 1 +ln x find du/dx

OpenStudy (anonymous):

1/x

OpenStudy (baru):

yes now multiply dy/du and du/dx

OpenStudy (anonymous):

you mean (1+2e^u)*(1/x)?

OpenStudy (baru):

yes, now that is dy/dx

OpenStudy (baru):

\[\frac{dy}{dx}=(1+2e^u)\times \frac{1}{x}\]

OpenStudy (anonymous):

would it be (1/x)(2e^u/x) and now that they are over the same denomentator then 2e^u

OpenStudy (baru):

i dont know what you mean... but have you followed till dy/dx ?

OpenStudy (anonymous):

yes (1 * 2e^u) * 1/x

OpenStudy (anonymous):

now i can distribute

OpenStudy (anonymous):

right?

OpenStudy (baru):

now, you need to plug these in \[x=1/e \\ u=1+\ln(1/e)\] you can simplify u a bit further before you plug in

OpenStudy (anonymous):

i got \[(1+2e^(1+\ln(1/x)))(1/(1/e))\]

zepdrix (zepdrix):

You have to put curly braces around the stuff you want in exponents :) e^{1+ln(1/x)} Sall good though, LaTeX takes a little getting used to ^^

OpenStudy (anonymous):

oh thank you

OpenStudy (baru):

i dont follow what you did, how do you still have an 'x' in there?

OpenStudy (anonymous):

\[(1+2e^{1+\ln(1/(1/e))})(1/(1/e))\]

OpenStudy (baru):

have a look at this \[u=1 + \ln(1/e) \\ u=1+\ln(e^{-1})\\u=1-\ln (e)\\u=1-1\\u=0\]

OpenStudy (anonymous):

oh goodness i didn't think of that

OpenStudy (baru):

\[\frac{dy}{dx}=(1+2e^u)\times \frac{1}{x} \\ plug ~ In\\ x=1/e ~and ~u=0\]

OpenStudy (anonymous):

2e

OpenStudy (baru):

check again... i get 3e

OpenStudy (anonymous):

oh i see how becasue 2e^0 is 2(e^0) and so it is 2(1)

OpenStudy (baru):

yep :)

OpenStudy (anonymous):

thank you so much

OpenStudy (baru):

sure :)

OpenStudy (anonymous):

sorry it took so long

OpenStudy (baru):

not at all :p

OpenStudy (baru):

@zepdrix anything to add?

zepdrix (zepdrix):

Nahhh c: gj guysss

OpenStudy (baru):

cool :P cuz u were making me a bit insecure xD

OpenStudy (anonymous):

no way you were doing a great job